OpenStudy (anonymous):

i have the equation 3x-2y=-16 i already put it in y=mx+b format but i don't know how to get the slope intercept

OpenStudy (anonymous):

slope (m) is the number in front of the x and the y-intercept is the other number (b).

OpenStudy (anonymous):

y = mx + b y = (3/2)x +8 Slope is 3/2 Y-intercept is 8 or (0,8)

OpenStudy (katrinakaif):

Slope is the number that comes before teh x, which in y= mx +b is m.. The y intercept is 'b'

OpenStudy (anonymous):

ok thanks